# F-Fidelity: A Robust Framework for Faithfulness Evaluation of Explainable AI

### v1

@@ 4 Empirical Verification of Robustness
To demonstrate the robustness of our proposed F-Fidelity framework, we conduct comprehensive experiments across multiple domains, including image classification, time series analysis, and natural language processing. Our evaluation strategy builds upon the concept introduced by Rong et al. (2022), which posits that an ideal evaluation method should yield consistent rankings in both MoRF and LeRF settings. We introduce a novel approach for a fair comparison using a degradation operation on a good explanation, such as an explanation obtained by Integrated Gradients (IG) (Sundararajan et al., 2017), generating a series of explanations with varying levels of random noise. This method provides a clear ground truth (GT) ranking based on noise level, allowing for controlled evaluation.
@@
We evaluate the performance of F-Fidelity against established baselines such as Fidelity (Fidelity), ROAR (Hooker et al., 2019), and R-Fidelity (Zheng et al., 2023) across a wide range of sparsity levels, from 5% to 95% at 5% intervals. Throughout our evaluation, we focus on three key Spearman rank correlations: “MoRF vs. GT”, “LeRF vs. GT”, and “MoRF vs. LeRF”. It allows us to assess the methods’ performance under various conditions, from highly sparse to nearly complete explanations. To provide a thorough analysis, we employ both macro and micro correlation metrics:
@@
  • Macro Correlation (Rong et al., 2022; Pan et al., 2021; Zhu et al., 2024): We compute the AUC with respect to sparsity across the entire 5-95% range for each explanation method. The macro correlations are then calculated using these AUC values, providing an overall performance measure across all sparsity levels.
@@ 4.1 Image Classification Explanation Evaluation
Setup. We use Cifar-100 (Krizhevsky et al., 2009) and Tiny-Imagenet (Deng et al., 2009) [^1] https://github.com/rmccorm4/Tiny-Imagenet-200?tab=readme-ov-file, as the benchmark datasets. To obtain a pre-trained model to be explained, we adopt ResNet (He et al., 2016). More experiments with Vision Transformer(ViT) (Dosovitskiy et al., 2020) can be found in Appendix D.4. To generate different explanations, we first use two explanation methods, SmoothGrad Squared (SG-SQ) (Smilkov et al., 2017) and GradCAM (Selvaraju et al., 2017) to obtain the explanations. Then we use a set proportion of noise perturbations in the image, $\left[0.0,0.2,0.4,0.6,0.8,1.0\right]$. We provided the implementation detail in Appendix C.
@@ 4.2 Time Series Classification Explanation Evaluation
Setup. We use two benchmark datasets for time series analysis: PAM for human activity recognition and Boiler for mechanical fault detection (Queen, 2023). For PAM, we use 534 samples across 8 activity classes, with each sample recorded using a fixed segment window length of 600 from 17 sensors. The Boiler dataset, used for mechanical fault detection, consists of 400 samples with 20 dimensions and a fixed segment window length of 36. We employ IG from the Captum library  [^2] https://github.com/pytorch/captum to obtain initial explanations. To generate different explanations, we apply noise perturbations to the importance of each timestamp, using proportions of $[0.0,0.1,0.2,0.3,0.4,0.5]$ for PAM and $[0.0,0.2,0.4,0.6,0.8,1.0]$ for Boiler. For R-Fidelity and F-Fidelity, we set $\alpha^{+}=\alpha^{-}=0.5$. Other settings remain consistent with the image classification task to ensure comparability across domains.
@@
Table 3: Spearman rank correlations and ranks on time series datasets. The best performance is marked as bold. the “-” means the correlation can’t be obtained because of the same rank of different explanations.
@@
5 Determining the Explanation Size via Fidelity Metrics
In many machine learning tasks, such as image classification and NLP, the ground truth explanations can be discretized into distinct clusters representing different levels of importance. For instance, in image classification, pixels associated with the target object tend to receive high importance scores. Conversely, pixels corresponding to the background or irrelevant regions receive low scores. However, in many practical scenarios, even good explainers that produce accurate explanation masks — as measured by the Fid and RFid evaluation metrics — may yield explanation scores that are not discretized into distinct clusters. For instance, when using PGExplainer on the MUTAG dataset (Luo et al., 2020), the explainer achieves over 90% accuracy in correctly ranking the influential edges of the input graph. Yet, the explanation scores are not clustered around discrete values. As a result, the size of the ground-truth explanation cannot be inferred from the explainer’s output since it provides only a ranking of edge importance without a clear separation into important and non-important edges.
In this section, we theoretically demonstrate that our proposed evaluation metric can recover the cluster sizes given an explainer that outputs the correct explanation mask (i.e., correctly ranks the importance of input elements). Thus, provided the explainer outputs an accurate mask function, our metric can recover the explanation size (also known as sparsity). To provide a concrete theoretical analysis, we first consider a classification problem under a set of idealized assumptions that generalize the above observations on the clustering of explanation scores. In the subsequent sections, we provide empirical evaluations demonstrating the applicability of our theoretical results to real-world settings.
Specifically, let us consider a classification task defined by a joint distribution $P_{\mathbf{X},Y}$ and a classifier $f:\mathbf{x}\mapsto y$. We assume that the input elements can be partitioned into several influence tiers. That is, for any given input $\mathbf{x}$, there exists a partition $\mathcal{C}_{k}(\mathbf{x}),k\in[r]$ of the index set $[t]\times[d]$, where $\mathcal{C}_{k}(\mathbf{x})$ represents the set of indices of the input elements belonging to tier $k$, and $c_{k}=|\mathcal{C}_{k}(\mathbf{x})|$ are the (fixed) tier sizes. For a given mask $\mathbf{m}$, the probability of correct classification based on the masked input $\mathbf{x}\odot\mathbf{m}$ depends only on the counts of unmasked elements in each influence tier. Formally,
$P(Y|\mathbf{x}\odot\mathbf{m})=g(j_{1},j_{2},\dots,j_{r}),$
where $g:[c_{1}]\times[c_{2}]\times\cdots\times[c_{r}]\to[0,1]$ is a function monotonically increasing with respect to the lexicographic ordering on its input, and $j_{k}\in[c_{k}]$ is the number of elements in $\mathcal{C}_{k}(\mathbf{x})$ whose corresponding mask element in $\mathbf{m}$ is non-zero (unmasked). We further focus on Shapley-value-based explanations, which provide a theoretical foundation for our analysis. Recall that the Shapley value associated with an element $(i,j)\in[t]\times[d]$ of $\mathbf{x}$ is given as (Lundberg, 2017):
$S_{\mathbf{x}}(i,j)=\sum_{\mathbf{m}:m_{i,j}=0}\frac{\|m\|_{1}!(td-\|m\|_{1}-1)!}{(td)!}(P(\mathbf{x}\odot\mathbf{m}^{\prime})-P(\mathbf{x}\odot\mathbf{m})),$
Theorem 1.
The proof is provided in the Appendix.
This theorem shows that $FFid^{+}$ can recover the size of the most influential tier (i.e., the first cluster size) when the explainer’s ranking is close to that of an ideal Shapley-based explainer. Specifically, the value of $s$ in which $FFid^{+}$ changes direction corresponds to the size of an influence tier. This is verified in the empirical evaluation provided in Section 6. This result implies that even when an explainer provides continuous scores without distinct clustering, our metric can infer the underlying discrete structure of the ground truth explanations and recover the explanation size.
6 Emprical Verification of determining the Explanation Size
In the terminology of Section 5, the pixels can be divided into two tiers: those belonging to the digit, which form the explanation, and those belonging to the background, which do not. The explanation size is thus $c_{1}=\gamma td$. According to Theorem 1, we expect $FFid^{+}$ to increase for $s<c_{1}$ and decrease for $s>\max\left(\frac{\beta}{\alpha^{+}}td,\gamma td\right)$. This behavior is illustrated in Figure 1 for various values of $\beta$ and $\gamma$. Furthermore, when $\frac{\beta}{\alpha^{+}}td<c_{1}$ (e.g., the blue curve in each plot), the point at which $FFid^{+}$ changes direction corresponds to the size of the explanation, $\gamma$. This allows us to recover $\gamma$ by evaluating $FFid^{+}$ over different $\beta$ values, as predicted by Theorem 1.
(a) $\gamma=0.1,\alpha^{+}=0.5$
(b) $\gamma=0.15,\alpha^{+}=0.5$
(c) $\gamma=0.2,\alpha^{+}=0.5$
(d) $\gamma=0.25,\alpha^{+}=0.5$
Figure 1: The relationship between ground truth explanation size $\gamma$ and $FFid^{+}$ in F-Fidelity.
@@ 8 Conclusion
In this paper, we introduced F-Fidelity, a robust framework for faithfulness evaluation in explainable AI. By leveraging a novel fine-tuning process, our method significantly mitigates the OOD problem that has plagued previous evaluation metrics. Through comprehensive experiments across multiple data modalities, we demonstrated that F-Fidelity consistently outperforms existing baselines in assessing the quality of explanations. Notably, our framework revealed a relationship between evaluation performance and ground truth explanation size under certain conditions, providing valuable insights into the nature of model explanations.
@@ Appendix A Detailed Algorithm
Algorithm 1 Computating $FFid^{+}$,$FFid^{-}$
1:  Input: model to be explained $f$, train set $\mathcal{D}_{t}$, explanation test set $\mathcal{D}_{v}$, explainer $\psi$, explanation size $s$, hyperparameters $\alpha^{+}$, $\alpha^{-}$, $\beta$, training epochs $E$, sampling numbers $T$,
2:  Output: $FFid^{+}$, $FFid^{-}$.
@@
5:     for  $(\mathbf{x}_{i},y)$ in $\mathcal{D}_{t}$ do
6:        update $f^{r}$ by using $\mathcal{L}\left(f\left(\mathbf{x}-P_{\beta}\odot\mathbf{x}\right),y\right)$
@@
11:     for  $t$ in range($T$do
12:        $\alpha^{+}=\text{min}\left(\alpha^{+},\frac{\beta}{s}\right)$
13:        $\alpha^{-}=\text{min}\left(\alpha^{-},\frac{\beta}{s}\right)$
14:        $FFid^{+}[i,t]\leftarrow\mathds{1}(y_{i}=f^{r}(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f^{r}(\chi^{+}(\mathbf{x}_{i},\alpha^{+},s))$
15:        $FFid^{-}[i,t]\leftarrow\mathds{1}(y_{i}=f^{r}(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f^{r}(\chi^{-}(\mathbf{x}_{i},\alpha^{-},s))$
@@
17:     $FFid^{+}[i]\leftarrow\frac{1}{T}\sum_{t=1}^{T}FFid^{+}[i,t]$
18:     $FFid^{-}[i]\leftarrow\frac{1}{T}\sum_{t=1}^{T}FFid^{-}[i,t]$
@@
20:  $FFid^{+}\leftarrow\frac{1}{|D_{v}|}\sum_{i=1}^{|D_{v}|}FFid^{+}[i]$
21:  $FFid^{-}\leftarrow\frac{1}{|D_{v}|}\sum_{i=1}^{|D_{v}|}FFid^{-}[i]$

### v2

@@ 5 Experiments
To demonstrate the robustness of the F-Fidelity framework, we conduct comprehensive experiments across multiple domains, including image classification, time series analysis, and natural language processing. Our evaluation strategy builds upon the concept introduced by Rong et al. (2022), which posits that an ideal evaluation method should yield consistent rankings in both MoRF and LeRF settings. To further establish a controlled experimental setting with ground truth (GT) rankings, we introduce a novel approach for a fair comparison using a degradation operation on a good explanation, such as an explanation obtained by Integrated Gradients (IG) (Sundararajan et al., 2017), generating a series of explanations with varying levels of random noise. Specifically, we first obtain initial (good) explanations using well-established explainers, then systematically degrade them by adding different ratios of random noise. Since explanation quality naturally decreases with increased noise, this creates a ground truth ranking where explanations with less noise should rank higher.
@@
We evaluate the performance of F-Fidelity against established baselines such as (Fidelity), ROAR (Hooker et al., 2019), and R-Fidelity (Zheng et al., 2024) across a wide range of sparsity levels, from 5% to 95% at 5% intervals. Throughout our evaluation, we focus on three key Spearman rank correlations that measure how well different evaluation metrics align with ground truth and each other. Specifically, “MoRF vs. GT” measures how well $Fid^{+}$ rankings align with ground truth rankings when removing important features first, while “LeRF vs. GT” measures the correlation between $Fid^{-}$ rankings and ground truth when retaining important features. The “MoRF vs. LeRF” correlation assesses the consistency between $Fid^{+}$ and $Fid^{-}$ evaluations, where strong negative correlation indicates that features identified as important by one metric are consistently identified as important by the other. These correlations allow us to assess the methods’ performance under various conditions, from highly sparse to nearly complete explanations. To provide a thorough analysis, we employ both macro and micro correlation metrics:
@@
  • Macro Correlation: Following Rong et al. (2022)’s approach of evaluating overall explainer consistency, and inspired by the AUC-based aggregation methods in Zhu et al. (2024) and Pan et al. (2021), we compute the AUC with respect to sparsity across the entire 5-95% range for each explanation method. The macro correlations are then calculated using these AUC values, providing an overall performance measure across all sparsity levels.
@@ 5.1 Image Classification Explanation Evaluation
Setup. We use CIFAR-100 (Krizhevsky et al., 2009) and Tiny-Imagenet (Deng et al., 2009) [^1] https://github.com/rmccorm4/Tiny-Imagenet-200?tab=readme-ov-file, as the benchmark datasets. To obtain a pre-trained model to be explained, we adopt ResNet (He et al., 2016). More experiments with Vision Transformer(ViT) (Dosovitskiy et al., 2020) can be found in Appendix D.1. To generate different explanations, we first use two explanation methods, SmoothGrad Squared (SG-SQ) (Smilkov et al., 2017) and GradCAM (Selvaraju et al., 2017) to obtain the explanations. Then we use a set proportion of noise perturbations in the image, $\left[0.0,0.2,0.4,0.6,0.8,1.0\right]$. We provide the implementation detail in Appendix C.
@@ 5.2 Time Series Classification Explanation Evaluation
Setup. We use two benchmark datasets for time series analysis: PAM for human activity recognition and Boiler for mechanical fault detection (Queen, 2023). For PAM, we use 534 samples across 8 activity classes, with each sample recorded using a fixed segment window length of 600 from 17 sensors. The Boiler dataset, used for mechanical fault detection, consists of 400 samples with 20 dimensions and a fixed segment window length of 36. We employ IG from the Captum library  [^2] https://github.com/pytorch/captum to obtain initial explanations. To generate different explanations, we apply noise perturbations to the importance of each timestamp, using proportions of $[0.0,0.1,0.2,0.3,0.4,0.5]$ for PAM and $[0.0,0.2,0.4,0.6,0.8,1.0]$ for Boiler.
@@
Table 3: Spearman ranks correlations and ranks on time series datasets with LSTM as classifier. The best performance is marked as bold. the “-” means the correlation can’t be obtained because of the same rank of different explanations.
@@
5.3 Natural Language Classification Explanation Evaluation
Setup. We use two benchmark datasets for our NLP experiments: the Stanford Sentiment Treebank (SST2) (Socher et al., 2013) for binary sentiment classification and the Boolean Questions (BoolQ) (Socher et al., 2013) dataset for question-answering tasks. For SST2, we utilize 67,349 sentences for training and 872 for testing. BoolQ comprises 9,427 question-answer pairs for training and 3,270 for testing. We employ two popular model architectures: LSTM networks and Transformer-based models (Appendix D.1). We follow the setting in image classification. Specifically, to generate explanations, we first use IG to obtain initial explanations, then apply noise perturbations to the importance of each timestamp at levels of $[0.0,0.2,0.4,0.6,0.8,1.0]$. We compare our F-Fidelity method against baselines including Fidelity and R-Fidelity, evaluating performance using both macro and micro Spearman correlations. We use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1e-4, keeping other hyperparameters at their default values.
Table 4: Spearman rank correlation and rank results with LSTM model on SST2 and BoolQ datasets.
Correlation SST2 BoolQ
Fidelity R-Fidelity F-Fidelity Fidelity R-Fidelity F-Fidelity
Macro Corr. MoRF vs GT $\downarrow$ -1.$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -1.$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -0.99$\pm$0.03 -1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.00
LeRF vs GT $\uparrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.01 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -0.99$\pm$0.02 -1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.00
Micro Rank MoRF vs GT $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.05$\pm$0.22 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
LeRF vs GT $\downarrow$ 1.05$\pm$0.22 1.05$\pm$0.22 1.05$\pm$0.22 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.11$\pm$0.31 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
Results. The results are shown in Table 3. Our experiments in the NLP domain reveal an interesting phenomenon that NLP models demonstrate remarkable robustness to OOD inputs, which is reflected in the performance of various fidelity metrics. Notably, the vanilla Fidelity (Fidelity) method achieves excellent results, often matching or closely approaching the performance of more sophisticated methods like R-Fidelity and our proposed F-Fidelity. This is evident from the near-perfect correlations (-1.00 or 1.00) across most metrics in both datasets. This strong performance of vanilla fidelity suggests that the OOD problem, which often necessitates more complex evaluation frameworks in other domains, may be less pronounced in NLP tasks. The robustness of NLP models to perturbations in input tokens likely contributes to this phenomenon, allowing simpler evaluation methods to maintain their efficacy. However, it’s worth noting that our proposed F-Fidelity method still demonstrates consistent top-tier performance across all metrics and models, reinforcing its versatility and reliability even in scenarios where simpler methods perform well.
5.4 Practical Guidelines for F-Fidelity Usage
We provide in-depth analysis on the relation between model robustness and performance of F-Fidelity in Appendix D.2. The observed robustness patterns exhibit a direct correlation with F-Fidelity’s effectiveness, which provides practical guidance. For tasks where models exhibit significant sensitivity to perturbations, such as computer vision and time series models, F-Fidelity should be the preferred choice over traditional fidelity metrics to ensure reliable explanation evaluation. In these domains, the substantial performance degradation under perturbation indicates vulnerability to OOD issues, making F-Fidelity’s robustness-enhancing properties particularly valuable. However, for tasks where models demonstrate inherent robustness to feature removal, such as many NLP tasks, simpler fidelity metrics are preferred. This relationship between model robustness and evaluation method choice enables one to make informed decisions based on their specific domain characteristics and accuracy requirements. In Appendix D.2, we provide a simple robustness test by measuring model accuracy under different masking ratios to determine the most appropriate evaluation metric.
@@ 7 Conclusion
In this paper, we introduced F-Fidelity, a robust framework for faithfulness evaluation in explainable AI. By leveraging a novel fine-tuning process, our method significantly mitigates the OOD problem that has plagued previous evaluation metrics. Through comprehensive experiments across multiple data modalities, we demonstrated that F-Fidelity consistently outperforms existing baselines in assessing the quality of explanations. Notably, our framework revealed a relationship between evaluation performance and ground truth explanation size under certain conditions, providing valuable insights into the nature of model explanations. In the future, we plan to explore alternative perturbation strategies, such as Gaussian blur. Additionally, we will investigate fine-tuning models with limited training data to further enhance the applicability and practicality of F-Fidelity, particularly in resource-constrained settings.
Acknowledgments
This project was partially supported by NSF grants IIS-2331908 and CCF-2241057. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Detailed Algorithm
Algorithm 1 Computing $FFid^{+}$,$FFid^{-}$
1:  Input: model to be explained $f$, train set $\mathcal{D}_{t}$, explanation test set $\mathcal{D}_{v}$ with $n=|\mathcal{D}_{v}|$ samples, explainer $\psi$, explanation size $s$, original parameters $\alpha^{+}_{orig}$, $\alpha^{-}_{orig}$, fixed fraction $\beta$, training epochs $E$, sampling numbers $N$
2:  Output: $FFid^{+}$, $FFid^{-}$
@@
5:     for $(\mathbf{x}_{i},y_{i})$ in $\mathcal{D}_{t}$ do
6:        update $f^{r}$ by using $\mathcal{L}(f(\mathbf{x}_{i}-P_{\beta}\odot\mathbf{x}_{i}),y_{i})$
@@
11:     for $k$ in range($N$do
12:        $\alpha^{+}=\min(\alpha^{+}_{orig},\frac{\beta td}{s})$
13:        $\alpha^{-}=\min(\alpha^{-}_{orig},\frac{\beta td}{td-s})$
14:        $FFid^{+}[i,k]\leftarrow\mathds{1}(y_{i}=f^{r}(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f^{r}(\chi^{+}(\mathbf{x}_{i},\alpha^{+},s)))$
15:        $FFid^{-}[i,k]\leftarrow\mathds{1}(y_{i}=f^{r}(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f^{r}(\chi^{-}(\mathbf{x}_{i},\alpha^{-},s)))$
@@
17:     $FFid^{+}[i]\leftarrow\frac{1}{N}\sum_{k=1}^{N}FFid^{+}[i,k]$
18:     $FFid^{-}[i]\leftarrow\frac{1}{N}\sum_{k=1}^{N}FFid^{-}[i,k]$
@@
20:  $FFid^{+}\leftarrow\frac{1}{|\mathcal{D}_{v}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{v}}FFid^{+}[i]$
21:  $FFid^{-}\leftarrow\frac{1}{|\mathcal{D}_{v}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{v}}FFid^{-}[i]$
@@ Appendix C Detailed Experiment Setup
In F-Fidelity, the hyperparameter $\beta$ is used in both fine-tuning and evaluation. We select $0.1$ by default. For the sampling number, $N$, a larger number helps reduce the standard deviation but leads to increasing evaluation time. In this paper, we select a balance point of $50$. The hyperparameter $\alpha^{+}$ and $\alpha^{-}$ are designed to alleviate the OOD problem, we choose $\alpha=\alpha^{+}=\alpha^{-}=0.5$. We provide hyperparameter sensitivity studies in Section D.9 and show comprehensive results in Section E.4. For the key parameter $\beta$, we further provide an in-depth analysis in Section D.3. For baselines, we use their default values as suggested in their original papers or codes.